OpenStudy (mysticfury):

WILL MEDAL this is my last question someone help What is the slope-intercept form of the function that contains the point (–1, 5) and has a slope of –4?

rebeccaxhawaii (rebeccaxhawaii):

i just told you

OpenStudy (boldjon):

y=mx+b y=-4x+b Plug in (-1,5) and solve for b And then just plug back in b

OpenStudy (boldjon):

Slope intercept is y= mx + b. With B being the y intercept. If it contains 1,1 and the slope is positive 6, do the equation 1 = 6 (1) + b. 1 = 6 + b -5 = b. So the answer is y = 6x - 5

OpenStudy (mysticfury):

i just wanna make sure @rebeccaxhawaii cause someone else had a different answer

rebeccaxhawaii (rebeccaxhawaii):

no the slope is not 6

rebeccaxhawaii (rebeccaxhawaii):

the slope is already given which is -4

OpenStudy (mysticfury):

nice pic @weasel123 but this time i do :D